OpenStudy (anonymous):

factor each expression. (each factor should have integer coefficients. 1600x^2+880xy+121y^2

OpenStudy (raden):

do completing square, that's your answer 1600x^2+880xy+121y^2 = (40x)^2 + 2(40x)(11y) + (11y)^2 = (40x + 11y)^2 = (40x + 11y)(40x + 11y)